# Selective Attention Improves Transformer

### v1

@@ Abstract
Unneeded elements in the attention’s context degrade performance. We introduce Selective Attention, a simple parameter-free change to the standard attention mechanism which reduces attention to unneeded elements. Selective attention improves language modeling performance in a variety of model sizes and context lengths. For example, a range of transformers trained with the language modeling objective on C4 with selective attention perform equivalently to standard transformers with $\sim$2X more heads and parameters in their attention modules. Selective attention also allows decreasing the size of the attention’s context buffer, leading to meaningful reductions in the memory and compute requirements during inference. For example, transformers with 100M parameters trained on C4 with context sizes of 512, 1,024, and 2,048 need 16X, 25X, and 47X less memory for their attention module, respectively, when equipped with selective attention, as those without selective attention, with the same validation perplexity.
@@ 1 Introduction
Figure 1: A visualization of the soft-masking by selective attention (red strike-through) and attention strength (averaged across heads, blue highlight) for different tasks (see Section 2).
@@ 2 Motivating Examples
Finally, Figure 1 also shows that for the trivial task of Parity*, where intermediate results are stored every other token, so that the model’s output is only a function of the last two tokens, everything but the last two tokens is masked. For the Copy task, selective attention persists the entirety of the string to be copied until copying starts, and then masks out every element as it is copied.
@@ 3 Selective Attention
Selective attention is a simple modification on top of standard attention. For context size $N$, we produce a real-valued $N\times N$ soft-mask matrix $S$ where $S_{i,j}$ denotes how much token $x_{i}$ masks token $x_{j}$ (see Section 3.1). We then constrain $S$, e.g. to be causal and non-negative (see Section 3.2). We finally accumulate the information in matrix $S$ into a new matrix $F$, taking into account masking by all previous tokens (see Section 3.3). The matrix $F=\text{Accumulate}(\text{Constrain}(S))$ is then simply subtracted from the attention logits before applying the softmax:
@@ 3.3 Accumulation
Figure 2: A sketch implementation of selective attention. The colored lines are the additions to standard attention. Note that selective attention adds a negligible amount of computation, however care should be taken in practice in order to not interfere with compiler optimizations.
@@ 4 Context Pruning
To see how, note that once a token is sufficiently masked by selective attention, it will not contribute meaningfully to any future attention operations. Such tokens can be safely evicted from the context buffer. We could pick a fixed threshold, and prune all elements whose soft masking is higher than the threshold (i.e. $F_{i,j}>\tau$), but that would make the memory savings hard to take advantage of (e.g. due to fragmentation and a variable number of dropped tokens each iteration). Instead, we observe that the sparsity (i.e. the magnitude of the soft-masking) per layer is stable across samples (see Section 7). To that effect we set a fixed memory budget for each layer, which directly translates to memory and compute savings. Since when a token is dropped it remains dropped for all future tokens, given a memory budget of $K=K_{1},\dots,K_{L}$ tokens for each layer, when processing the first $K_{l}$ tokens in layer $l$ we drop nothing, and for each following token we drop the single not-yet-dropped past token with the highest $F$ value. This maintains no more than $K_{l}$ tokens in layer $l$. Given an overall memory budget, we allocate it between the layers, using a greedy iterative approach. For context size $N$, we initialize $K^{0}_{1}=N,\dots,K^{0}_{L}=N$. In each iteration $t$, we set $K^{t}_{n}=K^{t-1}_{n}$ for all $n\neq m$, and $K^{t}_{m}=K^{t-1}_{m}-C$ for a constant $C$ (we use 8 in our experiments), where $m=\operatorname{argmin}_{i}PPL(\cdot|K^{t-1}-C_{i})$, where $C_{i}=(0,\dots,0,C,0,\dots,0)$ contains $C$ at the $i$th position. In other words, we iteratively reduce the memory budget of the layer that impacts model performance the least. We stop when model performance reaches a predefined threshold, in our experiments, the performance of a standard transformer without selective attention.
@@
$\mathcal{L}_{mem}=\mathcal{L}_{ppl}+\epsilon\cdot\frac{\sum_{l=1}^{L}\max_{i}M^{l}_{i}}{{L\cdot n_{\neq pad}}}$ (2)
Where $\mathcal{L}_{ppl}$ is the standard log-perplexity loss, $\epsilon$ is a small weight factor (in our experiments we set $\epsilon=0.1$ without further tuning), $L$ is the number of layers, $n_{\neq pad}$ is the number of non-pad tokens, and $M^{l}_{i}=i-\sum_{k=1}^{i}\min(F^{l}_{i,k},\tau)/\tau$ is our approximation for the memory requirements at the $i$th token for layer $l$ ($0\leq M^{l}_{i}\leq i$). We clamp $F^{l}_{i,k}$ from above by $\tau$ so as not to reward increasing it indefinitely ($F$ is already clamped from below to 0). We set $\tau=1$ without further tuning. Since the memory required for a given layer is the maximum memory required for each of the tokens, the loss only considers the maximum among the $M_{i}^{l}$s. We observe further reduction in context sizes with this explicit loss term (see Section 6.2).
@@ 6.1 Generation Quality
Transformers with selective attention perform consistently better, as measured by perplexity on the validation set, across model and context sizes. We also observe consistent improvement on a downstream task, HellaSwag (Zellers et al., 2019).
@@
Figure 4: Perplexity on the C4 validation set after 524,288 training steps of transformers with various sizes (parameterized by $d$ as per Section 5) with and without selective attention. For the cases without selective attention we add additional attention heads with their respective parameters (i.e. we increase the sizes of all projection matrices). Transformers with selective attention perform equivalently to those with standard attention modules with $\sim$2X as many heads and parameters.
In addition to perplexity on the validation set, we also measure model accuracy on the HellaSwag benchmark (Zellers et al., 2019), with and without selective attention, for models of various sizes (Table 1). We observe consistent improvements across all model sizes with selective attention.
Table 1: Comparing the accuracy of transformers of different sizes with and without selective attention on the HellaSwag benchmark.
Accuracy (baseline) Accuracy (selective attention)
$d=16$ 38.8% 40.0%
$d=18$ 41.8% 42.6%
$d=20$ 44.0% 45.1%
$d=22$ 46.0% 47.7%
$d=24$ 48.7% 50.3%
$d=26$ 51.0% 52.2%
$d=28$ 53.5% 53.8%
@@ 6.2 Inference Efficiency
For example, for $d=12$ transformers with context sizes of 512, 1,024, and 2,048, we see that with selective attention we can maintain the baseline’s perplexity while reducing the memory requirements of the attention module by factors of 5X, 7X, and 8X respectively, without any explicit losses (i.e. using only the standard language modeling objective). When training with the $\mathcal{L}_{mem}$ loss (Equation 2) and an $\epsilon$ value of 0.1, the improvements grow to 16X, 25X, and 47X respectively. We also measure the memory savings when considering only very long examples by filtering C4 to only include examples that are at least 90% the size of the context buffer. In that settings we get memory savings of 12X, 18X, and 24X, while maintaining the perplexity of the baseline without selective attention. Finally, to maintain the perplexity gains of selective attention, instead of matching the perplexity of the baseline (i.e. taking the rightmost point on the flat part of the blue graphs of Figure 6), we need 3X, 4X and 4X less memory, for the context sizes above. In all cases we optimized the per-layer budgets on a training set, while the reported results are on a separate unseen test set.
@@ 7 Selection Patterns
Figure 7 from Appendix A.8 illustrates the values of the $F$ matrix averaged over 1,000 examples, and demonstrates that the sparsity patterns are stable across examples. Interestingly, we observe that these sparsity patterns are sometimes stable across different training runs, hinting at some general properties of language modeling (see Figure 8 in Appendix A.8).
Figure 9 from Appendix A.8 depicts the items remaining in the context buffer after pruning (see Section 4), with a budget set to match the perplexity of a standard transformer. We observe that the per-layer budgets correspond well to the values of the $F$ matrix as seen in Figures 5 and 7. For example, layer 6 which gets the highest budget also has the lowest $F$ values. This might indicate that the scales of the $F$ values are consistent across layers.
@@ 8 Related Works
Context Pruning. A part of our work (Section 4) consists of removing elements from the context buffer. Several works employ this mechanism in order to increase inference efficiency. Among those are compression methods, that aim to learn a compressed representation for tokens in the context buffer, with or without auxiliary losses, in order to replace several tokens with their compressed form (Munkhdalai et al., 2024; Ren et al., 2023; Yun et al., 2023; Mu et al., 2024; Rae et al., 2019). Another line of work tries to simply remove elements from the context buffer without replacing them with new compressed forms, with minimal negative impact to model quality. The simplest and most widely adopted of these is using local attention windows in some of the layers (Wang et al., 2019; Beltagy et al., 2020; Zaheer et al., 2021). More sophisticated variants employ heuristics to evict less useful tokens from the context buffer instead of just the earliest ones (Oren et al., 2024; Zhang et al., 2023; Liu et al., 2023b; Berchansky et al., 2023; Ge et al., 2024). Dynamic context pruning (DCP) (Anagnostidis et al., 2024) proposed a mechanism for fine tuning existing models to make inference more efficient. They learn which tokens to prune from the context buffer via a mechanism which shares similarities with selective attention. DCP is more involved (e.g. needs root solving for evaluating the $\alpha$-sigmoid), introduces new parameters, requires an auxiliary loss, and produces binary prune decisions. Notably, DCP didn’t observe the perplexity gains we see from selective attention333 We only became aware of DCP after this work was concluded. It is interesting that both works share some important similarities, although developed from very different perspectives (see Section 10). .
Inference Efficiency. A part of our work (Section 4) consists of making inference from transformers more efficient. Many approaches aim to speed up inference from transformers, including distillation (Hinton et al., 2015), sparcification (Jaszczur et al., 2021), quantization (Hubara et al., 2016), architecture modification (So et al., 2022; Shazeer, 2019), and algorithmic optimization (Dao et al., 2022; Leviathan et al., 2022).
@@ 9 Discussion
Limitations and future directions. We applied selective attention to decoder-only transformers. It could be interesting to investigate its applicability to encoders as well; Reducing the size of the context as in Section 4 improves inference efficiency but not training efficiency. It might be interesting to explore iteratively reducing the size of the context buffer during training; We did not further train the models after removing elements as per Section 4. It seems conceivable that further improvements could be achieved with some additional training after context reduction; We only experimented with pre-training models with selective attention. It is interesting to investigate how it could be applied in a fine-tuning step to existing models; While we observed similar results with selective attention in several setups (Section 5), there are still important variants we did not test, notably transformers with multi-query (Shazeer, 2019) and grouped-query (Ainslie et al., 2023) attention, as well as models much larger than 1B parameters; Finally, it would be interesting to implement selective attention in a GPU-aware way, similar to Flash Attention (Dao et al., 2022).
@@ 10 Improving Neural Architectures
We are very strong advocates for this method, and believe that finding basic problems for which we cannot program a general solution by hand on a neural model, is an extremely fertile approach for producing further architecture improvements.
@@ A.1 Transformers with Selective Attention Learn a General Solution to Variable Assignment
A.2 Separate Bilinear Form
@@
A.3 Ablating the Constraints
@@
A.4 Self-Impact
$d=10$ $d=12$ $d=14$ $d=16$ $d=18$ $d=26$
Selective attention (no shift) 2.927 2.832 2.753 2.692 2.641 2.516
Selective attention 2.923 2.831 2.750 2.691 2.639 2.511
A.5 Perplexity-Efficiency Trade-Off
@@
A.6 Example Details
The following text from the C4 validation set was used in Figures 5, 9, and 10:

### v2

@@ Abstract
Unneeded elements in the attention’s context degrade performance. We introduce Selective Attention, a simple parameter-free change to the standard attention mechanism which reduces attention to unneeded elements. Selective attention consistently improves language modeling and downstream task performance in a variety of model sizes and context lengths. For example, transformers trained with the language modeling objective on C4 with selective attention perform language modeling equivalently to standard transformers with $\sim$2X more heads and parameters in their attention modules. Selective attention also allows decreasing the size of the attention’s context buffer, leading to meaningful reductions in the memory and compute requirements during inference. For example, transformers trained on C4 with context sizes of 512, 1,024, and 2,048 need 16X, 25X, and 47X less memory for their attention module, respectively, when equipped with selective attention, as those without selective attention, with the same validation perplexity.
@@ 1 Introduction
Figure 1: A visualization of the masking by selective attention (red strike-through) and attention strength (averaged across heads, blue highlight) for different tasks (see Section 2).
@@ 2 Motivating Examples
Figure 1 also shows that for the trivial task of Parity*, where intermediate results are stored every other token, so that the model’s output is only a function of the last two tokens, everything but the last two tokens is masked. For the Copy task, selective attention persists the entirety of the string to be copied until copying starts, and then masks out every element as it is copied. See Appendix A.2.
@@ 3 Selective Attention
Selective attention is a simple modification on top of standard attention. The key idea is that tokens can mask previous tokens, i.e. the amount of attention a token c pays a previous token a can be reduced by the tokens located between a and c. For context size $N$, we produce a real-valued $N\times N$ masking matrix $S$ where $S_{i,j}$ denotes how much token $x_{i}$ masks token $x_{j}$ (see Section 3.1). We then constrain $S$, e.g. to be causal and non-negative (see Section 3.2). We finally accumulate the information in matrix $S$ into a new matrix $F$, taking into account masking by all previous tokens (see Section 3.3). The matrix $F=\text{Accumulate}(\text{Constrain}(S))$ is then simply subtracted from the attention logits before applying the softmax:
@@ 3.3 Accumulation
Figure 2: A sketch implementation of selective attention. The colored lines are the additions to standard attention.
@@ 4 Context Pruning
To see how, note that once a token is sufficiently masked by selective attention, it will not contribute meaningfully to any future attention operations. Such tokens can be safely evicted from the context buffer. We could pick a fixed threshold, and prune all elements whose soft masking is higher than the threshold (i.e. $F_{i,j}>\tau$), but that would make the memory savings hard to take advantage of (e.g. due to fragmentation and a variable number of dropped tokens each iteration). Instead, we observe that the sparsity (i.e. the magnitude of the masking) per layer is stable across samples (see Section 7). To that effect we set a fixed memory budget for each layer, which directly translates to memory and compute savings. Since when a token is dropped it remains dropped for all future tokens, given a memory budget of $K=K_{1},\dots,K_{L}$ tokens for each layer, when processing the first $K_{l}$ tokens in layer $l$ we drop nothing, and for each following token we drop the single not-yet-dropped past token with the highest $F$ value. This maintains no more than $K_{l}$ tokens in layer $l$. Given an overall memory budget, we allocate it between the layers, using a greedy iterative approach. For context size $N$, we initialize $K^{0}_{1}=N,\dots,K^{0}_{L}=N$. In each iteration $t$, we set $K^{t}_{n}=K^{t-1}_{n}$ for all $n\neq m$, and $K^{t}_{m}=K^{t-1}_{m}-C$ for a constant $C$ (we use 8 in our experiments), where $m=\operatorname{argmin}_{i}\mathcal{L}(\cdot|K^{t-1}-C_{i})$, where $C_{i}=(0,\dots,0,C,0,\dots,0)$ contains $C$ at the $i$th position. In other words, we iteratively reduce the memory budget of the layer that impacts model performance the least. We stop when model performance reaches a predefined threshold, in our experiments, the performance of a standard transformer without selective attention.
@@
$\mathcal{L}_{mem}=\mathcal{L}+\epsilon\cdot\frac{\sum_{l=1}^{L}\max_{i}M^{l}_{i}}{{L\cdot n_{\neq pad}}}$ (2)
Where $\mathcal{L}$ is the standard loss (log-perplexity in the case of language modeling), $\epsilon$ is a small weight factor (in our experiments we set $\epsilon=0.1$ without further tuning), $L$ is the number of layers, $n_{\neq pad}$ is the number of non-pad tokens, and $M^{l}_{i}=i-\sum_{k=1}^{i}\min(F^{l}_{i,k},\tau)/\tau$ is our approximation for the memory requirements at the $i$th token for layer $l$ ($0\leq M^{l}_{i}\leq i$). We clamp $F^{l}_{i,k}$ from above by $\tau$ so as not to reward increasing it indefinitely ($F$ is already clamped from below to 0). We set $\tau=1$ without further tuning. Since the memory required for a given layer is the maximum memory required for each of the tokens, the loss only considers the maximum among the $M_{i}^{l}$s. We observe further reduction in context sizes with this explicit loss term (see Section 6.2).
@@ 6.1 Generation Quality
Transformers with selective attention perform consistently better, as measured by perplexity on the validation set, across model and context sizes. We also observe consistent improvements on a set of downstream tasks.
@@
Figure 4: Perplexity of transformers of various sizes with and without selective attention. For the cases without selective attention we add additional attention heads with their respective parameters (i.e. increase the sizes of all projection matrices). Transformers with selective attention perform equivalently to those with standard attention with $\sim$2X as many heads and parameters.
In addition to perplexity on the validation set, we also measure model accuracy on the ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2019), CommonSenseQA (Talmor et al., 2019), and OpenBookQA (Mihaylov et al., 2018) benchmarks, with and without selective attention, for models of various sizes (see Table 1). We observe consistent improvements across all model sizes with selective attention.
Table 1: Accuracy numbers for baseline transformers with standard attention and transformers with selective attention for various model sizes and downstream tasks.
ARC (Easy) ARC (Challenge) HellaSwag
Base Selective Base Selective Base Selective
$d=16$ 38.46% 39.33% 23.24% 23.24% 38.83% 40.01%
$d=18$ 40.48% 41.10% 24.67% 24.86% 41.77% 42.60%
$d=20$ 41.33% 41.85% 25.14% 25.52% 43.97% 45.08%
$d=22$ 42.43% 42.70% 25.83% 26.41% 46.04% 47.71%
$d=24$ 43.83% 44.47% 26.64% 26.76% 48.70% 50.32%
$d=26$ 44.70% 45.58% 26.37% 26.95% 51.02% 52.25%
$d=28$ 45.64% 46.95% 26.76% 27.37% 53.50% 53.76%
CommonSenseQA OpenBookQA PiQA
Base Selective Base Selective Base Selective
$d=16$ 25.44% 25.82% 34.26% 34.33% 68.07% 68.49%
$d=18$ 26.45% 26.69% 35.32% 35.42% 69.05% 69.88%
$d=20$ 26.30% 26.63% 35.42% 35.27% 69.82% 70.49%
$d=22$ 26.56% 27.10% 35.96% 36.88% 70.55% 71.43%
$d=24$ 27.49% 27.49% 36.36% 37.13% 71.34% 71.95%
$d=26$ 27.70% 27.78% 37.64% 37.52% 71.89% 72.47%
$d=28$ 27.86% 28.51% 37.33% 37.54% 72.32% 72.85%
@@ 6.2 Inference Efficiency
For example, for $d=12$ transformers with context sizes of 512, 1,024, and 2,048, we see that with selective attention we can maintain the baseline’s perplexity while reducing the memory requirements of the attention module by factors of 5X, 7X, and 8X respectively, without any explicit losses (i.e. using only the standard language modeling objective). When training with the $\mathcal{L}_{mem}$ loss (Equation 2) and an $\epsilon$ value of 0.1, the improvements grow to 16X, 25X, and 47X respectively. We also measure the memory savings when considering only very long examples by filtering C4 to only include examples that are at least 90% the size of the context buffer. In that settings we get memory savings of 12X, 18X, and 24X, while maintaining the perplexity of the baseline without selective attention. To maintain the perplexity gains of selective attention, instead of matching the perplexity of the baseline (i.e. the rightmost point on the flat part of the blue graphs of Figure 6), we need 3X, 4X and 4X less memory, for the context sizes above, respectively.
In all cases above, we optimized the per-layer budgets on a training set and reported results on a separate unseen test set.
Finally, we compare selective attention to other attention variants and pruning methods and observe that it achieves a substantially better quality-cost trade-off than all baselines (see Appendix A.10).
@@ 7 Selection Patterns
Figure 9 in Appendix A.12 illustrates the values of the $F$ matrix averaged over 1,000 examples, and demonstrates that the sparsity patterns are stable across examples. Interestingly, we observe that these sparsity patterns are sometimes stable across different training runs, hinting at some general properties of language modeling (see Figure 10 in Appendix A.12).
Figure 11 in Appendix A.12 depicts the items remaining in the context buffer after pruning (see Section 4), with a budget set to match the perplexity of a standard transformer. We observe that the per-layer budgets correspond well to the values of the $F$ matrix as seen in Figures 5 and 9. For example, layer 6 which gets the highest budget also has the lowest $F$ values. This might indicate that the scales of the $F$ values are consistent across layers.
@@ 8 Related Works
Context Pruning. A part of our work (Section 4) consists of removing elements from the context buffer. Several works employ this mechanism in order to increase inference efficiency. Among those are compression methods, that aim to learn a compressed representation for tokens in the context buffer, with or without auxiliary losses, in order to replace several tokens with their compressed form (Munkhdalai et al., 2024; Ren et al., 2023; Yun et al., 2023; Mu et al., 2024; Rae et al., 2019). Another line of work tries to simply remove elements from the context buffer without replacing them with new compressed forms, with minimal negative impact to model quality. The simplest and most widely adopted of these is using local attention windows in some of the layers (Wang et al., 2019; Beltagy et al., 2020; Zaheer et al., 2021). More sophisticated variants employ heuristics to evict less useful tokens from the context buffer instead of just the earliest ones (Oren et al., 2024; Zhang et al., 2023; Liu et al., 2023b; Berchansky et al., 2023; Ge et al., 2024). Anagnostidis et al. (2024) fine tunes existing models to learn to prune tokens from the context buffer with similarities to selective attention, but is more involved (e.g. needs root solving for evaluating the $\alpha$-sigmoid), produces binary prune decisions, and notably doesn’t improve quality.
Inference Efficiency. A part of our work (Section 4) consists of making inference from transformers more efficient. Many approaches aim to speed up inference from transformers, including distillation (Hinton et al., 2015), sparsification (Jaszczur et al., 2021), quantization (Hubara et al., 2016), architecture modification (So et al., 2022; Shazeer, 2019), and algorithmic optimization (Dao et al., 2022; Leviathan et al., 2022).
@@ 9 Discussion
Future directions. We applied selective attention to decoder-only transformers. It could be interesting to investigate its applicability to encoders as well (see Appendix A.9 for initial results in an encoder-decoder setup); Reducing the size of the context as in Section 4 improves inference efficiency but not training efficiency. It might be interesting to explore iteratively reducing the size of the context buffer during training; We did not further train the models after removing elements as per Section 4. It seems conceivable that further improvements could be achieved with some additional training after context reduction; We only experimented with pre-training models with selective attention. It is interesting to investigate how it could be applied in a fine-tuning step to existing models; While we observed similar results with selective attention in several setups (Section 5), there are still important variants we did not test, notably transformers with multi-query (Shazeer, 2019) and grouped-query (Ainslie et al., 2023) attention, as well as models much larger than 1B parameters; Finally, selective attention can be implemented in a GPU-aware way, similar to Flash Attention (Dao et al., 2022).
@@ 10 Improving Neural Architectures
We are strong advocates of this method, and believe that finding basic problems for which we cannot program a general solution by hand on a neural model, is a fertile approach for architecture improvements.
@@ A.1 Transformers with Selective Attention Learn a General Solution to Variable Assignment
See Appendix A.11 for an example of the attention patterns for the Variable Assignment task.
A.2 Additional Synthetic Tasks
We validate selective attention on two additional synthetic tasks, Copy and Parity, that are on the two opposite extremes in terms of memory requirements. See Figure 1.
Setup. We train small transformers ($d=3$), with and without selective attention, to solve the Copy and Parity tasks. We train with a batch size of 2,048 for 65,536 steps.
Copy. Here the transformer gets an arbitrary sequence of varying length delimited by a special token and is tasked with outputting a copy of the sequence. We used lengths that are uniformly distributed between 1 and 24. The context size for this task is $3+2\times L_{max}$, where $L_{max}=24$ is the length of the longest possible input sequence (the 3 extra tokens are for <BOS>, <EOS> and the special end-of-input-sequence token). To solve this task, the model cannot forget anything before copying starts, after which point it can mask out tokens that were already copied.
Parity. Here the transformer gets a binary sequence where bits in the odd positions are random and bits in the even positions contain the parity of all bits in the earlier odd positions. Equivalently, bits in the even positions contain the XOR of the bits in the previous two positions, so the two previous positions are enough for computing the next token. The loss only considers the even positions.
Unsurprisingly, transformers with selective attention achieve practically 0 loss (and 100% accuracy) at the end of training, as do standard transformers without selective attention, on both tasks.
A.3 Separate Bilinear Form
@@
A.4 Self-Impact
$d=10$ $d=12$ $d=14$ $d=16$ $d=18$ $d=26$
Selective attention (no shift) 2.927 2.832 2.753 2.692 2.641 2.516
Selective attention 2.923 2.831 2.750 2.691 2.639 2.511
A.5 Ablating the Constraints
@@
A.6 Perplexity-Efficiency Trade-Off
@@
A.7 Example Details
The following text from the C4 validation set was used in Figures 5, 11, and 12: